# Consistency conditions for O-plane unsmearing from second-order perturbation theory

### v1

@@ 3.3 Near-source behavior of leading corrections
The equations (3.20) fully determine the dominant behavior of the leading order corrections near the source loci, even without having to find a global solution. As this behavior will play a very important role in the second order calculation, we write it here explicitly. Near any one particular source, transverse spherical symmetry is restored to leading order in the transverse distance. We can use local coordinates $y^{m}$, such that the metric is diagonal at the source locus and $y^{1,2,3}$ are longitudinal to the source, while $y^{4,5,6}$ are transverse. We can also define a radial coordinate $r$ that measures proper transverse distance from the source, defined by
$dr^{2}=g^{(0)\perp}_{mn}dy^{m}dy^{n}=g^{(0)}_{44}dy^{4}dy^{4}+g^{(0)}_{55}dy^{5}dy^{5}+g^{(0)}_{66}dy^{6}dy^{6}\ .$ (3.24)
@@
$\Box\Phi=-\frac{\mu}{\tau^{(0)}}\frac{\delta(r)}{r^{2}}+...$ (3.25)
@@
$g^{(1)\|}_{mn}=\frac{\mu}{2}g^{(0)}_{mn}\frac{1}{r}+...$ $=\frac{1}{2}g^{(0)}_{mn}\tau^{(0)}\Phi+...$ (3.26)
@@
which can be checked to satisfy the equations (3.20) near the source. Note in particular that
$g^{(1)m}_{~{}~{}m}=0+...$ (3.27)
@@
$F_{2}^{(1)}=-\mu\sin\theta d\theta\wedge d\phi+...\ ,$ (3.28)
@@
$F_{mn}=-\mu\ \epsilon^{\perp}_{mnp}\frac{y^{p}}{r^{3/2}}+...\ ,$ (3.29)
@@
$|F_{2}|^{2}=\frac{\mu^{2}}{r^{4}}+...=\tau^{(0)2}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(i)}+...\ .$ (3.30)
@@ 4.1 Second order warp factor equation
00 0 $=\tau^{(0)2}\frac{\Box w^{(2)}}{w^{(0)}}+\mathcal{E}_{w,\text{sing.}}+\mathcal{E}_{w,\text{reg.}}\ ,$ (4.1)
@@
$\mathcal{E}_{w,\text{sing.}}$ $=-\frac{1}{2}|F_{2}^{(1)}|^{2}-2\frac{\tau^{(0)2}}{w^{(0)2}}w^{(1)}\Box w^{(1)}+6\frac{\tau^{(0)2}}{w^{(0)2}}\partial_{m}w^{(1)}\partial^{m}w^{(1)}$ (4.2)
$\quad+4\frac{\tau^{(0)}}{w^{(0)}}\tau^{(1)}\Box w^{(1)}+4\frac{\tau^{(0)}}{w^{(0)}}\partial_{m}\tau^{(1)}\partial^{m}\tau^{(1)}$
@@
$\mathcal{E}_{w,\text{reg.}}$ $=R^{(AdS_{4})}\frac{\tau^{(0)2}}{w^{(0)2}}\left(\frac{w^{(1)}}{w^{(0)}}-\frac{\tau^{(1)}}{\tau^{(0)}}\right)$ (4.3)
@@
To obtain a constraint on the first order corrections, we proceed similarly to the example of Einstein gravity on a torus from section 2. The idea is to integrate equation (4.1) over the internal manifold. This integral must, of course, vanish. However, unlike the linearized gravity example, it’s not obvious that this integration eliminates the $\Box w^{(2)}$ term, despite its appearance as a total derivative. Indeed, the equation (4.1) must be satisfied everywhere on the internal manifold, but some terms have singular near-source behavior. A priori, it may be the case that some of this divergence is cancelled by the $\Box w^{(2)}$ term. This would mean that the integral of $\Box w^{(2)}$ alone over the full manifold is ill-defined and integrating the full equation does not produce a constraint on the first-order corrections alone. Excising small neighborhoods around the source loci from the domain of integration doesn’t help, since $w^{(2)}$ would then contribute to the integral via boundary terms at the excisions that wouldn’t vanish even as the size of the excised region goes to zero.
@@
$\frac{1}{\tau^{(0)2}}\mathcal{E}_{w,\text{sing.}}=-\frac{5}{4}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(i)}-\frac{5}{4}\Phi^{(i)}\Box\Phi^{(i)}+...\ ,$ (4.4)
where, again the ... denotes terms that have at most $1/r$ divergence. The expression (4.4) appears to be a total derivative, but both terms are singular near the source. The first term has $r^{-4}$ behavior, and thus must be canceled by the $\Box w^{(2)}$ term in (4.1). This results in
$w^{(2)}=\mu^{2}\frac{w^{(0)}}{\tau^{(0)2}}\frac{5}{16r^{4}}+...\ ,\qquad r>0\ .$ (4.5)
@@
$\frac{\delta(r)}{r^{3}}\ ,$ (4.6)
which takes us outside the realm of integrable functions (and distributions as limits thereof) and into more exotic generalized functions. This is a common occurence in non-linear theories and gravity in particular. A natural intepretation of such terms is that they determine the boundary conditions on the boundary of a small excised ball around the singularity, in the limit of vanishing ball radius, in such a way that integration by parts works. [^3] More rigorous treatments of such objects exist in the literature e.g. [42] in the context of general relativity. It would be interesting to apply these methods here, but for our purposes this simple treatment will suffice. Thus, $w^{(2)}$ is given by (4.5) for $r>0$, but must also contain a singular generalized function piece at $r=0$, which specifically cancels the divergence of its integral. Any remaining finite contributions to $\Box w^{(2)}$ form a regular total derivative whose integral is well defined over the entire compact manifold and therefore vanishes. This allows us to proceed with the argument in the style of section 2 and write down the following integral constraint on the first order corrections.
$\int d^{6}y\sqrt{g^{(0)}}\left(\mathcal{E}_{w,\text{sing.}}+\mathcal{E}_{w,\text{reg.}}\right)=0\ .$ (4.7)
@@
  • The equality of the coefficients in equation (4.4) is not trivial since (4.2) is not a manifest total derivative. For example, in (4.2), only the second and third lines take the form of total derivatives. The first line only does so because the near-source behavior of the flux has the appropriate coefficent compared to the gradient of the warp factor. This, in turn, is only true because the O-plane has equal charge and tension. The cancellation of the source term in the last line of (4.2) is also non-trivial and results from the identical near-source behavior of the dilaton and the trace of the transverse metric.
  • The $-5/4$ coefficient in (4.4) is precisely the one required to reproduce the expansion of the warp factor for a spherically symmetric O6 solution with flat worldvolume
    $w=(c-\mu/r+\mathcal{O}(r^{2}))^{-1/4}\ ,$ (4.8)
@@
    The agreement with the behavior of the massless O6-plane might appear to suggest that the near-source behavior of the full solution is different from the proposed numerical solutions of [43] for a localized O6-plane in massive type IIA and more similar to those in the massless theory. However, although at this order the $1/r^{2}$ terms match up, we expect further $1/n$ corrections to appear to this term if the computation is pushed to higher order. Thus it is difficult to foresee what the final function will resum to and whether it may indicate instead a behavior more similar to the solutions of [43].
  • We have so far not discussed the effect of possible O-plane intersections. Indeed, terms in (4.2) and (4.14) also contain cross-terms of the form
    $\Phi^{(i)}\Box\Phi^{(j)}\ ,\qquad\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(j)}$ (4.9)
@@
    $\frac{1}{r_{1}^{2}}\frac{\delta(r_{2})}{r_{2}^{2}}\ .$ (4.10)
    For an intersection such that no longitudinal directions are shared, performing the integral over the directions transverse to the delta function we are left with a remaining integrand proportional to $1/r_{1}^{2}$ with the coefficient fixed by the intersection angles. Integrated over the remaining directions, this only gives a finite contribution. The same is true for the $\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(j)}$ terms, which are proportional to
    $\frac{1}{r_{1}}\frac{1}{r_{2}}.$ (4.11)
    For sources with no common directions, each factor is integrated over a separate subspace and gives a finite result.
@@
    $\frac{1}{2}F^{(i)}_{mn}F^{(j)mn}=\tau^{(0)2}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(j)}\ ,$ (4.12)
    which precisely completes the near-intersection behavior of the first line of (4.2) to a total derivative.
    Thus intersections do not make any additional singular contributions to the integral constraints. At most, they give a finite contribution to the integration constants in the first-order corrections.
    Note that the necessary cancellations between the flux and other terms would not hold true for an intersection with a shared longitudinal direction. Fortunately, such O-plane intersections (including the ones resulting from the orientifold involution used in the original DGKT construction) are inconsistent, as explained in [44]. On the other hand, the allowed type of intersection can be eliminated by blow-ups, in which case no extra divergent contributions to the integral constraint are possible. It is thus perhaps not that surprising that they are also absent in the intersecting limit.
@@ 4.2 Second order dilaton equation, other constraints and overall consistency
00 0 $=\tau^{(0)}\Box\tau^{(2)}+\mathcal{E}_{\tau,\text{sing.}}+\mathcal{E}_{\tau,\text{reg.}}\ ,$ (4.13)
@@
$\mathcal{E}_{\tau,\text{sing.}}$ $=\frac{3}{2}|F_{2}^{(1)}|^{2}+8\frac{\tau^{(0)}}{w^{(0)}}\partial_{m}\tau^{(1)}\partial^{m}w^{(1)}$ (4.14)
$\quad+2\tau^{(1)}\Box\tau^{(1)}+2\partial_{m}\tau^{(1)}\partial^{m}\tau^{(1)}-2g^{(1)}_{mn}\nabla^{m}\partial^{n}\tau^{(1)}-2\nabla^{m}g^{(1)}_{mn}\partial^{n}\tau^{(1)}$
@@
$\mathcal{E}_{\tau,\text{reg.}}$ $=-\frac{3}{4}F^{(0)}_{mn}F^{(2)mn}-\frac{3}{2}g^{(1)mn}|F_{2}^{(0)}|^{2}_{mn}$ (4.15)
@@
Substituting the near-source expressions into $\mathcal{E}_{\tau,\text{sing.}}$ gives the near-source expression
$\frac{1}{\tau^{(0)2}}\mathcal{E}_{\tau,\text{sing.}}=\frac{3}{4}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(i)}+\frac{3}{4}\Phi^{(i)}\Box\Phi^{(i)}+...\ .$ (4.16)
@@
$\tau^{(2)}=-\mu^{2}\frac{3}{16r^{4}}+...\ ,\qquad r>0\ .$ (4.17)
This is precisely the $1/r^{4}$ term in the large-$c$ expansion of
$(c-\mu/r+\mathcal{O}(r^{2}))^{3/4}\ .$ (4.18)
@@
$\int d^{6}y\sqrt{g^{(0)}}\left(\mathcal{E}_{\tau,\text{sing.}}+\mathcal{E}_{\tau,\text{reg.}}\right)=0\ ,$ (4.19)
@@
Finally, we might consider other constraints coming from the internal metric equation. One obvious and universally present constraint arises from the trace of the internal equation. The second-order expression is rather cumbersome, due to the multitude of possible index contractions, so we simply quote the results.
The final expression for the singular terms in the near-source limit:
$\frac{1}{2}\tau^{(0)2}\nabla^{m}\nabla^{n}g^{(2)}_{mn}-\frac{1}{2}\tau^{(0)2}\Box g^{(2)m}_{~{}~{}~{}m}$ $-\tau^{(0)}\Box\tau^{(2)}-2\frac{\tau^{(0)2}}{w^{(0)}}\Box w^{(2)}$ (4.20)
$=\tau^{(0)2}\left(\frac{21}{8}\partial_{m}\Phi\partial^{m}\Phi+\frac{21}{8}\Phi\Box\Phi\right)+...\ .$
@@
$g_{mn}=(c-\mu/r+\mathcal{O}(r^{2}))^{\pm 1/2}\ ,$ (4.21)
with $+$ in the transverse directions and -- for the longitudinal ones.

### v2

@@ 3.3 Near-source behavior of leading corrections
The equations (31) fully determine the dominant behavior of the leading order corrections near the source loci, even without having to find a global solution. As this behavior will play a very important role in the second order calculation, we write it here explicitly. Near any one particular source, transverse spherical symmetry is restored to leading order in the transverse distance. We can use local coordinates $y^{m}$, such that the metric is diagonal at the source locus and $y^{1,2,3}$ are longitudinal to the source, while $y^{4,5,6}$ are transverse. We can also define a radial coordinate $r$ that measures proper transverse distance from the source, defined by
$dr^{2}=g^{(0)\perp}_{mn}dy^{m}dy^{n}=g^{(0)}_{44}dy^{4}dy^{4}+g^{(0)}_{55}dy^{5}dy^{5}+g^{(0)}_{66}dy^{6}dy^{6}\ .$ (35)
@@
$\Box\Phi=-\frac{\mu}{\tau^{(0)}}\frac{\delta(r)}{r^{2}}+...$ (36)
@@
$g^{(1)\|}_{mn}=\frac{\mu}{2}g^{(0)}_{mn}\frac{1}{r}+...$ $=\frac{1}{2}g^{(0)}_{mn}\tau^{(0)}\Phi+...$ (37)
@@
which can be checked to satisfy the equations (31) near the source. Note in particular that
$g^{(1)m}_{~{}~{}m}=0+...$ (38)
@@
$F_{2}^{(1)}=-\mu\sin\theta d\theta\wedge d\phi+...\ ,$ (39)
@@
$F_{mn}=-\mu\ \epsilon^{\perp}_{mnp}\frac{y^{p}}{r^{3/2}}+...\ ,$ (40)
@@
$|F_{2}|^{2}=\frac{\mu^{2}}{r^{4}}+...=\tau^{(0)2}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(i)}+...\ .$ (41)
@@ 4.1 Second order warp factor equation
00 0 $=\tau^{(0)2}\frac{\Box w^{(2)}}{w^{(0)}}+\mathcal{E}_{w,\text{sing.}}+\mathcal{E}_{w,\text{reg.}}\ ,$ (42)
@@
$\mathcal{E}_{w,\text{sing.}}$ $=-\frac{1}{2}|F_{2}^{(1)}|^{2}-2\frac{\tau^{(0)2}}{w^{(0)2}}w^{(1)}\Box w^{(1)}+6\frac{\tau^{(0)2}}{w^{(0)2}}\partial_{m}w^{(1)}\partial^{m}w^{(1)}$ (43)
$\quad+4\frac{\tau^{(0)}}{w^{(0)}}\tau^{(1)}\Box w^{(1)}+4\frac{\tau^{(0)}}{w^{(0)}}\partial_{m}\tau^{(1)}\partial^{m}w^{(1)}$
@@
$\mathcal{E}_{w,\text{reg.}}$ $=R^{(AdS_{4})}\frac{\tau^{(0)2}}{w^{(0)2}}\left(\frac{w^{(1)}}{w^{(0)}}-\frac{\tau^{(1)}}{\tau^{(0)}}\right)$ (44)
@@
To obtain a constraint on the first order corrections, we proceed similarly to the example of Einstein gravity on a torus from section 2. The idea is to integrate equation (42) over the internal manifold. This integral must, of course, vanish. However, unlike the linearized gravity example, it’s not obvious that this integration eliminates the $\Box w^{(2)}$ term, despite its appearance as a total derivative. Indeed, the equation (42) must be satisfied everywhere on the internal manifold, but some terms have singular near-source behavior. A priori, it may be the case that some of this divergence is cancelled by the $\Box w^{(2)}$ term. This would mean that the integral of $\Box w^{(2)}$ alone over the full manifold is ill-defined and integrating the full equation does not produce a constraint on the first-order corrections alone. Excising small neighborhoods around the source loci from the domain of integration doesn’t help, since $w^{(2)}$ would then contribute to the integral via boundary terms at the excisions that wouldn’t vanish even as the size of the excised region goes to zero.
@@
$\frac{1}{\tau^{(0)2}}\mathcal{E}_{w,\text{sing.}}=-\frac{5}{8}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(i)}-\frac{5}{8}\Phi^{(i)}\Box\Phi^{(i)}+...\ ,$ (45)
where, again the ... denotes terms that have at most $1/r$ divergence. The expression (45) appears to be a total derivative, but both terms are singular near the source. The first term has $r^{-4}$ behavior, and thus must be canceled by the $\Box w^{(2)}$ term in (42). This results in
$w^{(2)}=\mu^{2}w^{(0)}\frac{5}{16r^{2}}+...\ ,\qquad r>0\ .$ (46)
@@
$\frac{\delta(r)}{r^{3}}\ ,$ (47)
which takes us outside the realm of integrable functions (and distributions as limits thereof) and into more exotic generalized functions. This is a common occurence in non-linear theories and gravity in particular. A natural intepretation of such terms is that they determine the boundary conditions on the boundary of a small excised ball around the singularity, in the limit of vanishing ball radius, in such a way that integration by parts works. [^3] More rigorous treatments of such objects exist in the literature e.g. Steinbauer:2006qi in the context of general relativity. It would be interesting to apply these methods here, but for our purposes this simple treatment will suffice. Thus, $w^{(2)}$ is given by (46) for $r>0$, but must also contain a singular generalized function piece at $r=0$, which specifically cancels the divergent part of its integral. Any remaining finite contributions to $\Box w^{(2)}$ form a regular total derivative whose integral is well defined over the entire compact manifold and therefore vanishes. This allows us to proceed with the argument in the style of section 2 and write down the following integral constraint on the first order corrections.
$\int d^{6}y\sqrt{g^{(0)}}\left(\mathcal{E}_{w,\text{sing.}}+\mathcal{E}_{w,\text{reg.}}\right)=0\ .$ (48)
@@
  • The equality of the coefficients in equation (45) is not trivial since (43) is not a manifest total derivative. For example, in (43), only the second and third lines take the form of total derivatives. The first line only does so because the near-source behavior of the flux has the appropriate coefficent compared to the gradient of the warp factor. This, in turn, is only true because the O-plane has equal charge and tension. The cancellation of the source term in the last line of (43) is also non-trivial and results from the identical near-source behavior of the dilaton and the trace of the transverse metric.
  • The $-5/8$ coefficient in (45) is precisely the one required to reproduce the expansion of the warp factor for a spherically symmetric O6 solution with flat worldvolume
    $w=(c-\mu/r+\mathcal{O}(r^{2}))^{-1/4}\ ,$ (49)
@@
    The agreement with the behavior of the massless O6-plane might appear to suggest that the near-source behavior of the full solution is different from the proposed numerical solutions of Saracco:2012wc for a localized O6-plane in massive type IIA and more similar to those in the massless theory. However, although at this order the coefficients of the $1/r^{2}$ terms match up, we expect further $1/n$ corrections to appear to this term if the computation is pushed to higher order. Thus it is difficult to foresee what the final function will resum to and whether it may indicate instead a behavior more similar to the solutions of Saracco:2012wc .
  • We have so far not discussed the effect of possible O-plane intersections. Indeed, terms in (43) also contain cross-terms of the form
    $\Phi^{(i)}\Box\Phi^{(j)}\ ,\qquad\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(j)}$ (50)
@@
    $\frac{1}{r_{1}}\frac{\delta(r_{2})}{r_{2}^{2}}\ .$ (51)
    For an intersection such that no transverse directions along the internal space are shared, performing the integral over the directions transverse to the delta function we are left with a remaining integrand proportional to $1/r_{1}$ with the coefficient fixed by the intersection angles. Integrated over the remaining directions, this only gives a finite contribution. The same is true for the $\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(j)}$ terms, which are proportional to
    $\frac{1}{r_{1}^{2}}\frac{1}{r_{2}^{2}}.$ (52)
    For a pair of sources with no common transverse directions, each factor is integrated over a separate subspace and gives a finite result.
@@
    $\frac{1}{2}F^{(i)}_{mn}F^{(j)mn}=\tau^{(0)2}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(j)}\ ,$ (53)
    which precisely completes the near-intersection behavior of the first line of (43) to a total derivative form. Thus these types of intersections do not make any additional singular contributions to the integral constraints. At most, they give a finite contribution to the integration constants in the first-order corrections.
    Note that the necessary cancellations between the flux and other terms, necessary for the near-source behavior to take on a “total derivative" form, would not hold true for an intersection with a shared transverse direction. Fortunately, such O-plane intersections in the DGKT construction result from an inconsistent choice of spacetime involution, as explained in Junghans:2023yue . On the other hand, the allowed type of intersection features no common transverse directions and can be eliminated by blow-ups, in which case no extra divergent contributions to the integral constraint are possible. It is thus perhaps not that surprising that they are also absent in the intersecting limit.
  • Our analysis is restricted to the two-derivative equations of motion. As emphasized in Andriot:2023fss , higher derivative terms begin to contribute at the same order as the second order perturbative corrections. One may therefore ask what their effect on our analysis might be. Since these terms involve higher derivatives of the first-order corrections, their near-source behavior will be more singular than any of the two-derivative terms.
    From here there are several possibilities. The best case scenario is that the higher-derivative terms appear with the appropriate coefficients so that their singular parts once again assemble into a “total derivative"-like structure involving a divergent function in the neighborhood of the O-plane locus and a generalized function at the locus itself such that the total integral is finite. The effect of the higher derivative terms would then be to correct the consistency conditions through finite contributions, which are highly suppressed away from the O-plane locus.
    A more puzzling alternative would be that the higher derivative terms introduce mismatched singular contributions, invalidating integration by parts when integrating over the manifold and forbidding the formulation of integral constraints altogether. This would also complicate the interpretation of the generalized function terms localized at the O-plane locus, since they would no longer be equivalent to setting boundary conditions for small excisions.
    Although the first option would certainly be more appealing for our analysis, without knowing all the higher derivative corrections that contribute at a given order, it is impossible to say which option plays out. That said, it is still a relevant question whether two-derivative solutions exist to begin with, before worrying how they might be corrected or invalidated by higher-derivative corrections. The perturbative unsmearing procedure promises a systematic approximation scheme to a two-derivative solution. Our present analysis suggests that this procedure remains consistent at higher orders when applied to the second-order equations of motion.
@@ 4.2 Second order dilaton equation, other constraints and overall consistency
00 0 $=\tau^{(0)}\Box\tau^{(2)}+\mathcal{E}_{\tau,\text{sing.}}+\mathcal{E}_{\tau,\text{reg.}}\ ,$ (54)
@@
$\mathcal{E}_{\tau,\text{sing.}}$ $=\frac{3}{2}|F_{2}^{(1)}|^{2}+8\frac{\tau^{(0)}}{w^{(0)}}\partial_{m}\tau^{(1)}\partial^{m}w^{(1)}$ (55)
$\quad+2\tau^{(1)}\Box\tau^{(1)}+2\partial_{m}\tau^{(1)}\partial^{m}\tau^{(1)}-2\tau^{(0)}g^{(1)}_{mn}\nabla^{m}\partial^{n}\tau^{(1)}-2\tau^{(0)}\nabla^{m}g^{(1)}_{mn}\partial^{n}\tau^{(1)}$
@@
$\mathcal{E}_{\tau,\text{reg.}}$ $=-\frac{3}{4}F^{(0)}_{mn}F^{(2)mn}-\frac{3}{2}g^{(1)mn}|F_{2}^{(0)}|^{2}_{mn}$ (56)
@@
Substituting the near-source expressions into $\mathcal{E}_{\tau,\text{sing.}}$ gives
$\frac{1}{\tau^{(0)2}}\mathcal{E}_{\tau,\text{sing.}}=\frac{3}{8}\partial_{m}\Phi^{(i)}\partial^{m}\Phi^{(i)}+\frac{3}{8}\Phi^{(i)}\Box\Phi^{(i)}+...\ .$ (57)
@@
$\tau^{(2)}=-\mu^{2}\frac{3}{16r^{2}}+...\ ,\qquad r>0\ .$ (58)
This is precisely the $1/r^{2}$ term in the large-$c$ expansion of
$(c-\mu/r+\mathcal{O}(r^{2}))^{3/4}\ .$ (59)
@@
$\int d^{6}y\sqrt{g^{(0)}}\left(\mathcal{E}_{\tau,\text{sing.}}+\mathcal{E}_{\tau,\text{reg.}}\right)=0\ ,$ (60)
@@
Finally, we might consider other constraints coming from the internal metric equation. One obvious and universally present constraint arises from the trace of the internal metric equation. The second-order expression is rather cumbersome, due to the multitude of possible index contractions, so we simply quote the results.
The final expression for the singular terms in the near-source limit is
$\frac{1}{2}\tau^{(0)2}\nabla^{m}\nabla^{n}g^{(2)}_{mn}-\frac{1}{2}\tau^{(0)2}\Box g^{(2)m}_{~{}~{}~{}m}$ $-\tau^{(0)}\Box\tau^{(2)}-2\frac{\tau^{(0)2}}{w^{(0)}}\Box w^{(2)}$ (61)
$=-\tau^{(0)2}\left(\frac{21}{8}\partial_{m}\Phi\partial^{m}\Phi+\frac{21}{8}\Phi\Box\Phi\right)+...\ .$
@@
$g_{mn}=(c-\mu/r+\mathcal{O}(r^{2}))^{\pm 1/2}\ ,$ (62)
with $+$ in the exponent for the transverse directions and -- for the longitudinal ones.